DX Field Data - 20140424 UT

Observers: JStorm, IIlyn (LBTO)
Telescope Operator: DHuerta (LBTO)
AO Support:: DMiller (Tucson)
Telescope Support:: JUrban (LBTO)
Instrument Support:: JMorris (LBTO)
Engineering Support: DAshby (Tucson)

Summary

Very good conditions, seeing around 0.8arcsec, low wind. Unfortunately we lost a lot of time to problems with the M2 on SX

We managed to verify hotspot positions and image quality on all beam splitters as well as measuring photometric zero points while having the pinhole guiding active all the time.

Details

9:00 The seeing blew up from 1 to more than 3arcsec in a few minutes and stayed up for maybe 5 minutes before falling back. Observed both on the DIMM and on the MODS guider.

9:10 Handover (2:10AM Local time) from OSU/RC.

9:30 Problems with M2 on DX...

9:35 Ilya finds that the status of the AGW filter wheels are not properly set. He always gets a return value of 0 even if the value should be between 1 and 3.

9:40 Problems with AGw 7+8 communications. Necessary to restart both.

9:55 M2 DX ripped again

10:20 Doug M. did a restart of M2 from Tucson and we restarted the pointing.

10:30 Doing spiral search on SX trying to find the star. It turned out to be that we had the ND2 filter in on the SX side.

10:32 Finally we have the right star and pointing on both eyes.

10:32 Presetting via PEPSI interface to HIP76984

10:33 Success.

10:35 Seeing about 0.9arcsec on both sides. It looks clear outside.

10:35 Testing the pinhole guiding, 100micron pinhole, BS1, ND0

10:40 After numerous WFS iterations the FWHM is down to 0.9 on SX and 0.7 arcsec on DX

10:50 Doing photometric ZP measurements and hotspot positions for the various beam splitters. Reference on BS1 from above.

10:53 Switch to BS2. Forgot to pause guiding.

10:56 New preset to HIP76984 with BS2

10:57 Guiding.

11:05 BS2 Hotspot SX: 366,378; DX: 332,374 Seeing: 0.73, 0.61arcsec

11:10 Changed to BS3 (no new preset, just paused guiding and changed BS)

11:14 BS3 Hotspot SX: 364,378; DX: 331,374 Seeing: 0.73, 0.61arcsec

11:15 Changed to BS4 (no new preset, just paused guiding and changed BS)

11:14 BS4 Hotspot SX: 372,379; DX: Seeing: 0.73, 0.61arcsec

11:18 M2 on DX ripped...

11:26 Called Doug M. to investigate M2 problem.

11:35 Renewed the preset to HIP76984 with BS4

11:40 BS4 Hotspot SX: 374,379; DX: 344,374 Seeing: 0.80, 1.5arcsec still not quite converged, but almost).

11:41 Switched to BS5

11:47 BS5 Hotspot SX: 372,378; DX: 340,373 Seeing: 0.80, 0.77

11:48 Switched to BS1

11:55 BS1 Hotspot SX: 368,376; DX: 339,373 Seeing: 0.73, 0.69 Airmass: 1.34 Ilya revised the code to update the guiding hotspot location to the value proper for the BS. This uses an absolute setting of guiding hotspot location using SetGuidingHotspot.

11:59 Checked BS2 and BS3 with revised SW, looks good.

12:03 Checking BS5, also looks fine!

12:06 Test with 200micron pinhole and BS5. We have a lot of back reflected light from the pinhole microlens.

12:11 Test with 200micron pinhole and BS1. Here we also have a lot of back reflected light from the pinhole microlens.

12:15 Switched to filter ND2. Warning appeared that guiding should be paused before changing filter wheel position. The filter DID change to the requested value (filter 1).

12:17 Paused guiding on SX and moved filter to 0 (ND0) The same warning appeared! The filter DID move.

12:18 Moved (without pausing) to filter 3 on SX. Same warning, but the filter DID move.

12:25 Stopped for twilight.
Topic revision: r1 - 25 Apr 2014, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback